1) the average of 3 test scors; if the first 2 scores are 78 and 82; s is the third score (how do you find the value of 's'?)
2) the sum of 3 consecutive integars, n=the 1st integr of the three (how do you find each numbe?)

Answer by stanbon(75887)

1) the average of 3 test scors; if the first 2 scores are 78 and 82; s is the third score (how do you find the value of 's'?)
---
average = (78+82+s)/3
-------------------------------
2) the sum of 3 consecutive integers, n=the 1st integr of the three (how do you find each numbe?)
sum = n+(n+1)+(n+2) = 3n+3
---------------------------
Cheers,
Stan H.